Circular No. 6189

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 O1

The following additional precise positions have been reported:

1995 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 24.59517 18^{h}43^{m}06.^{s}68 −32^{o}08′28′′.5  Abe
24.60142 18 43 06.47 −32 08 27.6  _{”}
24.85549 18 42 55.65 −32 08 03.7 12  Tichá
24.85676 18 42 55.61 −32 08 03.0  _{”}
24.85757 18 42 55.60 −32 08 03.2  _{”}
24.85836 18 42 55.54 −32 08 03.4  _{”}
24.85919 18 42 55.50 −32 08 03.0  _{”}
24.87581 18 42 54.81 −32 08 00.6  Pettarin
24.93351 18 42 52.37 −32 07 55.9  _{”}
25.23715 18 42 39.13 −32 07 25.0  Schleicher
25.24229 18 42 38.92 −32 07 24.8  _{”}
25.24516 18 42 38.82 −32 07 24.7  _{”}
25.52113 18 42 27.45 −32 06 58.5  Yoshida
25.52545 18 42 27.26 −32 06 57.9  _{”}
25.52975 18 42 27.07 −32 06 58.0 11.6  _{”}

H. Abe (Yatsuka). 0.26-m f/4.8 reflector + CCD. Communicated by S.

Nakano, Sumoto, Japan. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. E. Pettarin and A. Toso (Farra d’Isonzo). 0.4-m f/4.5 reflector + CCD. D. G. Schleicher and T. B. Spahr (Anderson Mesa Station, Lowell Obser-

vatory). 1.1-m f/8 Hall reflector + CCD. S. Yoshida (Wakayama). 0.25-m f/6.3 Schmidt-Cassegrain reflector + CCD.

Communicated by Nakano.

COMET 6P/D’ARREST

Total visual magnitude estimates (cf. IAUC 6179): July 3.40 UT, 12.2 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 4.40, 12.0 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 8.44, 11.5 (Hale); 9.30, 10.7 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 23.32, 8.8 (Morris, 20×80 binoculars; coma diameter ∼ 14′); 24.95, 9.4 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + V filter + CCD; coma diameter ∼ 15′).

1995 July 25 Daniel W. E. Green